OpenStudy (anonymous):

Katie has to wear a uniform to school. When she dresses each morning, she has a choice of three different shirt colors (white, green, or yellow) and two different skirt colors (green or plaid).

OpenStudy (kropot72):

Each of the 3 choices of shirt color can be combined with the 2 choices of skirt color. The number of outfits to choose from is \[3\times 2=?\]

OpenStudy (anonymous):

Sports Feet is having a sale on the Rebound basketball shoe. The shoe comes in four different colors (black, red, white, or gold) and two different materials (canvas or leather). How many different styles of this shoe are available?

OpenStudy (anonymous):

im really bad at problem solving=/

OpenStudy (kropot72):

For the question regarding shoe styles, use the same method that I showed you for the uniform question.

OpenStudy (anonymous):

i still dont understand

OpenStudy (anonymous):

nvm got it

OpenStudy (anonymous):

x - 3y = 4 2x - 6y = 8 how do i solve this by substitution